OpenStudy (wondermath):

Can anyone give me an example of a function that is CONTINUOUS but not DIFFERENTIABLE? (other than the absolute function) Thanks!

OpenStudy (amistre64):

piecewise functions

OpenStudy (wondermath):

but those arent continous tho

OpenStudy (jamesj):

How about this: f(x) = (x-1)^{1/3} It is everywhere continuous, but not differentiable at x = 1

OpenStudy (asnaseer):

what about this one: http://en.wikipedia.org/wiki/Weierstrass_function

OpenStudy (zarkon):

\[\sum_{k=0}^{\infty}a^k\cos(b^k\pi x)\] \[0<a<1;ab>1;b>1\]

OpenStudy (jamesj):

(Yes, the construction of everywhere continuous and nowhere differentiable functions is a famous topic in mathematics, but fairly tricky!)

OpenStudy (asnaseer):

More examples here: http://math.mit.edu/classes/18.013A/HTML/chapter06/section03.html

OpenStudy (zarkon):

Mathematical brownian motion

OpenStudy (asnaseer):

Zarkon - is that a description of the Weierstrass function?

OpenStudy (zarkon):

brownian motion ...no

OpenStudy (asnaseer):

I have never heard of mathematical brownian motion

OpenStudy (zarkon):

We should note that mathematical brownian motion is continuous and not differentiable with probability 1

OpenStudy (asnaseer):

I just looked it up - very interesting: http://en.wikipedia.org/wiki/Brownian_motion

OpenStudy (zarkon):

http://en.wikipedia.org/wiki/Wiener_process

OpenStudy (asnaseer):

yes - that is mentioned in the article I found as well.

OpenStudy (amistre64):

f(x) = x, for x <=1 = x^2 for x >1 is continuous ... and not diffied at x=1

OpenStudy (amistre64):

works better maybe at the x=0 junction, but same premise :)

OpenStudy (wondermath):

amistre, how come it's not differentiable?

OpenStudy (amistre64):

if we choose the: f(x) = x , x<=0 x^2, x>0 we have a cusp, a sharp, nonsmooth, junction at x=0. slope of the line from the left does not match the slope of the line from the right and therefore has no definiable slope for the point at x=0. For the same reason that |x| is not differentiable at x=0 ...

OpenStudy (amistre64):

in other words; the slope from the left is 1 ; the slope from the right is 0

OpenStudy (wondermath):

oh okayyyy thanks